# A Review of Graph-Powered Data Quality Applications for IoT Monitoring Sensor Networks

### v1

@@ Abstract
The development of Internet of Things (IoT) technologies has led to the widespread adoption of monitoring networks for a wide variety of applications, such as smart cities, environmental monitoring, and precision agriculture. A major research focus in recent years has been the development of graph-based techniques to improve the quality of data from sensor networks, a key aspect for the use of sensed data in decision-making processes, digital twins, and other applications. Emphasis has been placed on the development of machine learning and signal processing techniques over graphs, taking advantage of the benefits provided by the use of structured data through a graph topology. Many technologies such as the graph signal processing (GSP) or the successful graph neural networks (GNNs) have been used for data quality enhancement tasks. In this survey, we focus on graph-based models for data quality control in monitoring sensor networks. Furthermore, we delve into the technical details that are commonly leveraged for providing powerful graph-based solutions for data quality tasks in sensor networks, including missing value imputation, outlier detection, or virtual sensing. To conclude, we have identified future trends and challenges such as graph-based models for digital twins or model transferability and generalization.
@@ 1 Introduction
The GSP field emerged to translate the classic signal processing concepts, related to the analysis of time-dependent signals, to irregular domains such as graphs, defining the notion of signals defined over graphs [1, 2]. The use of graph-based models had already been a key element in applications such as route planning (e.g., Dijkstra’s algorithm) [3], community detection (e.g., clique percolation or Louvain algorithms) [4] or the analysis of complex networks such as biological and social networks [5, 6]. The representation of manifolds by means of graphs in the field of semi-supervised learning is another example of graph-powered applications [7]. Overall, the GSP framework [8] has enabled the use and development of novel techniques on data residing in graphs, thus emerging as an alternative to classical machine learning techniques that do not make explicit use of data structure. In this way, the graph topology, which represents the relationships between the graph’s nodes, is fed to graph-based models that explicitly model the structure of the data [9].
@@
A wide variety of concepts have been applied to signals defined over graphs, such as signal shift, translation, convolution, or filtering [10]. An important concept of the GSP is the notion of signal smoothness, also expressed via the total variation (TV) or the Dirichlet energy, which are quadratic forms and can be used to evaluate how a signal fits a given graph structure or vice-versa [11]. This idea is linked with the Graph Discrete Fourier Transform (GDFT) that makes use of the graph topology to obtain the graph Fourier basis and allow the computation of the transform coefficients of a graph signal and also led to the development of graph filters [12]. In the field of machine learning, graphs have been used as regularizers in optimization problems, e.g., the regularization of neural networks for semi-supervised learning tasks [13]. Moreover, the structured representation of the data allows for improving the interpretability of the models as well as the interpretability of their results [9]. Given the growing popularity of black-box models (e.g., deep neural networks), much emphasis has been placed on the interpretability of the techniques to facilitate their use. In addition, one of the greatest successes of graph-based techniques has been the development of GNNs, allowing the use of neural networks on data residing in irregular domains such as graphs, from more general message-passing-based GNNs to spectral networks based on the graph convolution operation [14, 15].
@@
Recently, graphs have unveiled their potential to represent Internet of Things (IoT) monitoring sensor networks, allowing the use of signal processing and machine learning techniques to carry out applications using sensor network data [16, 17]. Monitoring sensor networks are platforms whose purpose is to monitor the evolution of certain phenomena in a specific area of interest or scenario, e.g., precision agriculture systems [18], air quality monitoring [19], or industrial processes [20, 21]. These networks are of particular interest because they monitor critical phenomena or assets and provide data for further applications or decision-making processes. Figure 1 depicts different scenarios where IoT monitoring networks are used, common use cases include smart metering, industrial applications, and environmental monitoring among others. The applications aim to exploit the network measurements as well as their intrinsic structure to carry out several tasks. The quality of data from monitoring sensor networks is critical since the data can feed various applications or decision-making processes that require complete and accurate data. In the literature, data quality tasks have been developed to provide continuous, complete and accurate data. Some tasks include the detection of outliers and anomalies [22], the use of predictive models to create virtual sensors [23], the imputation of missing data [24], and the detection of clusters and communities among others [16]. Indeed, graph-based models provide alternatives to classic data-driven approaches and have been proven to be superior in terms of performance, possibility of distributed approaches, and interpretability. Henceforth, a deep understanding of signal processing and machine learning over graph approaches may unveil new applications and advantages for IoT monitoring sensor networks.
@@
Recent surveys are based on the technical review of models based mainly on GSP and GNN, mentioning some applications in areas such as computer vision or natural language processing [14, 25, 26, 27, 28]. Others focus on the applications of GNNs and graph-based models in areas such as the IoT, showing graph-based applications in cases such as environmental monitoring or biological data [17, 29, 30, 31]. In this survey, we focus on the application of graph-based models to the recently booming area of data quality improvement for monitoring IoT platforms [32, 31]. In particular, we not only cover GNNs applications but also provide complete coverage of graph-based modeling of monitoring networks, covering GSP, machine learning over graphs, and GNNs. We also present recent advances and applications in data quality control and provide a taxonomy for the most novel graph-based applications and trends.
@@
The outline of this paper is as follows; Section 2 motivates the content of the paper. Section 3 introduces technical concepts regarding signal processing and machine learning over graphs. Section 4 reviews different graph applications in IoT monitoring sensor networks, and section 5 presents future trends and challenges. Finally, section 6 concludes the manuscript. Figure 2 shows the outline of the survey.
@@ 2 Background and Motivation
Before the appearance of the GSP, graphs were already used in the field of graph-based semi-supervised learning to present the relationships of data residing in a manifold [25, 33]. In fact, label propagation emerged as a transductive method to estimate unlabeled data points from labeled ones, and inspired the GNN framework [34]. The advent of the GSP was a revolution since it allowed the translation of solid signal processing fundamentals to the realm of signals defined over graphs. Thus, this allowed the analysis of measurements residing on a graph via graph filters, signal shifts, signal convolutions, or the spectral graph Fourier transforms among others. In turn, the graph signal convolution led to the translation of convolutional neural networks to the graph paradigm, making use of the graph signal convolution operator, e.g., the Chebnet GNN [35]. Besides all the guarantees in terms of efficiency and accuracy, the GSP-based techniques provide interpretability, which has received a lot of attention due to the growing use of black box models, given the use of the graph shift operator, i.e., the matrix that describes the relationships between network nodes. Moreover, many of the techniques, e.g., graph filters, allow a distributed implementation, which is very interesting in case of resource limitations or in case of having nodes with computational resources, e.g., edge computing approach [36]. All this together makes the GSP provide appropriate tools to analyze the measurements produced by a network of sensors, being able to define these measurements as signals that reside in a graph.
@@
Machine learning over graphs usually refers to machine learning techniques that have been adapted to take into account that a graph can represent the data, or analogously that the data resides in a manifold. In this sense, graph-based semi-supervised learning is aimed at the application of algorithms to partially-labeled data residing on manifolds represented by graphs [7]. This idea has also been used in different works where the optimization function of different machine learning techniques has been modified to include a regularization term that takes into account that the data points reside on a graph [37]. Among some examples, we can find predictive models, e.g., kernel ridge regression [38] or regularized support vector regression [39], which can be very useful in the case of sensor network measurements. Even though, there exists a large range of tasks and applications that can be approached using machine learning, e.g., the target of the machine learning task can be estimating features at graph nodes, at graph edges, or at the graph as a whole. Throughout the manuscript, we restrict to the tasks that are useful in the sensor network domain.
@@
Although GNNs belong to the field of machine learning over graphs, in this paper we want to dedicate their own space to them due to their great importance and widespread adoption nowadays. Some well-known architectures include the GraphSAGE [40] or the Chebnet [35]. Indeed, the graph convolution operator allowed the application of the convolution neural networks to the graph realm, resulting in graph convolutional neural networks such as the Chebnet, in which the convolution is realized via a graph filter. Given the widespread application of neural networks, tailored GNN architectures are being continuously developed to tackle tasks such as missing value imputation, anomaly detection, data compression, event detection, or signal reconstruction [14]. Another example of disruption in this field is the emergence of graph attention networks and graph transformers, being the analogs of attention mechanisms and transformers in conventional neural networks [41].
@@
Figure 3 depicts the scope of this review. From now on we focus on GSP, machine learning over graphs, and GNNs, placing special emphasis on novel applications for data quality in monitoring sensor networks. Data quality is a key challenge in monitoring sensor networks, with tasks ranging from anomaly detection to missing value imputation [32]. Although existing surveys cover various topics related to GSP and GNN for IoT (Table 1), the review of graph-based models for data quality in IoT is not covered yet. To this end, this survey provides a novel description of graph-based models for data quality tasks in IoT monitoring sensor networks. In summary, our contributions are:
@@
    describe the foundations of different graph-based models (GSP, ML over graphs and GNNs) that can be used for improving data quality in IoT monitoring sensor networks,
@@
    provide a taxonomy of tasks that can help practitioners introduce graph-based models in monitoring sensor networks to improve data quality, and review recent work,
@@
    distinguish future trends and challenges that are emerging and that will gain interest in incoming years.
Table 1: Comparison between the scope of the manuscript with respect to other relevant surveys.
@@ 3 Graph-Based Modeling for Sensor Networks
In this section, we elaborate on the fundamentals of the GSP, machine learning over graphs, and GNNs, highlighting current trends and novel techniques. We denote the measurement of a given sensor $i$ at a given time $t$ as $x_{i}(t){\in}\mathbb{R}$, $N$ denotes the number of sensors, and $\mathbf{x}_{i}{\in}\mathbb{R}^{T}$ represents a vector of $T$ measurements recorded by the $i$-th sensor. Matrices are denoted by bold uppercase letters, vectors by bold lowercase letters, and sets by calligraphic letters.
@@
Figure 4: Example of how a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{S}\}$ can be used to represent a sensor network and the sensors’ relationships. $e_{ij}$ represents an edge connecting nodes $i$ and $j$. The entry of the graph shift matrix $S_{ij}$ assigns a weight to the edge $e_{ij}$, describing the strength of the relationship between nodes $i$ and $j$.
In addition, there are two variants of graphs that can be applied in sensor networks: i) static graphs $\mathcal{G}$ and ii) dynamic or time-varying graphs $\mathcal{G}^{(t)}$ whose structure depends on time $t$ [50]. Statics graphs can be used to represent sensor networks that are static or fixed so the relationships between the sensors are not expected to change. In contrast, dynamics graphs have a time-dependent structure, meaning that the relationships between the sensors as well as the set of sensors composing the graph can change over time. This approach is useful to represent mobile sensor networks or dynamic networks where sensors can join and leave the network and the relationship between them can change. We can further classify the graphs into three types; i) undirected, in which the relationships between nodes are reciprocal, i.e., $(i,j){\in}\mathcal{E}$ implies $(j,i){\in}\mathcal{E}$, or $S_{ij}{=}S_{ji}$, ii) directed, in which the relationships do not need to be reciprocal, and iii) multigraphs, in which more than one edge can exist for a given pair of nodes $(i,j)$. Figure 5 represents the different types of graphs. In the specific case of monitoring sensor networks, where a graph structure is used to represent the relationship between the sensors’ data, it makes sense to assume that the relationships between the sensor measurements are reciprocal, e.g., as in the case of correlation coefficients.
@@
Figure 5: Types of graphs: above, static graphics; below, dynamic graphics; $u$ denotes undirected, $d$ denotes directed, $m$ denotes multigraph, and $\mathcal{G}_{u}^{(t)}$ denotes a dynamic undirected graph at time $t$. $W_{ij}$ is the entry of the weight matrix that represents the edge $e_{ij}$ and $T{\in}\mathbb{N}$ represents a time offset.
@@
In addition, before moving on to explain the technical aspects of signal processing and machine learning over graphs, it is important to highlight the use of two types of predictive models in this graph-based approach; i) inductive models, and ii) transductive models. The objective of inductive learning is to infer a model from training instances, then this model can be applied to different testing instances. In contrast, in the case of transductive learning, training instances are used to obtain estimates for specific test instances. As we will see later, even though in the realm of sensor networks many models can be used in both ways, there are specific models of both types, such as a graph filter (inductive) or a graph kernel ridge regression (transductive). Figure 6 illustrates the idea behind these two learning paradigms.
@@
Figure 6: Illustration of inductive and transductive learning paradigms for graph-based models. In black training data, and in blue testing data. Inductive models are trained to perform inferences in other tests sets while transductive models are trained to perform inferences for a specific test set.

### v2

@@ Abstract
The development of Internet of Things (IoT) technologies has led to the widespread adoption of monitoring networks for a wide variety of applications, such as smart cities, environmental monitoring, and precision agriculture. A major research focus in recent years has been the development of graph-based techniques to improve the quality of data from sensor networks, a key aspect of the use of sensed data in decision-making processes, digital twins, and other applications. Emphasis has been placed on the development of machine learning (ML) and signal processing techniques over graphs, taking advantage of the benefits provided by the use of structured data through a graph topology. Many technologies such as graph signal processing (GSP) or the successful graph neural networks (GNNs) have been used for data quality enhancement tasks. This survey focuses on graph-based models for data quality control in monitoring sensor networks. In addition, it introduces the technical details that are commonly used to provide powerful graph-based solutions for data quality tasks in sensor networks, such as missing value imputation, outlier detection, or virtual sensing. To conclude, different challenges and emerging trends have been identified, e.g., graph-based models for digital twins or model transferability and generalization.
@@ 1 Introduction
The GSP field emerged to translate the classic signal processing concepts, related to the analysis of time-dependent signals, to irregular domains such as graphs, defining the notion of signals defined over graphs [1, 2]. The use of graph-based models had already been a key element in applications such as route planning (e.g., Dijkstra’s algorithm) [3], community detection (e.g., clique percolation or Louvain algorithms) [4] or the analysis of complex networks such as biological and social networks [5, 6]. The representation of manifolds by means of graphs in the field of semi-supervised learning is another example of graph-powered applications [7]. Overall, the GSP framework [8] has enabled the use and development of novel techniques on data residing in graphs, thus emerging as an alternative to classical ML techniques that do not make explicit use of data structure. In this way, the graph topology, which represents the relationships between the graph’s nodes, is fed to graph-based models that explicitly model the structure of the data [9].
A wide variety of concepts have been applied to signals defined over graphs, such as signal shift, translation, convolution, or filtering [10]. An important concept of the GSP is the notion of signal smoothness, also expressed via the total variation (TV) or the Dirichlet energy, which are quadratic forms and can be used to evaluate how a signal fits a given graph structure or vice-versa [11]. This idea is linked with the Graph Discrete Fourier Transform (GDFT) that makes use of the graph topology to obtain the graph Fourier basis and allow the computation of the transform coefficients of a graph signal and also led to the development of graph filters [12]. In the field of ML, graphs have been used as regularizers in optimization problems, e.g., the regularization of neural networks for semi-supervised learning tasks [13]. Moreover, the structured representation of the data allows for improving the interpretability of the models as well as the interpretability of their results [9]. Given the growing popularity of black-box models (e.g., deep neural networks), much emphasis has been placed on the interpretability of the techniques to facilitate their use. In addition, one of the greatest successes of graph-based techniques has been the development of GNNs, allowing the use of neural networks on data residing in irregular domains such as graphs, from more general message-passing-based GNNs to spectral networks based on the graph convolution operation [14, 15].
Recently, graphs have unveiled their potential to represent IoT monitoring sensor networks, allowing the use of signal processing and ML techniques to carry out applications using sensor network data [16, 17]. Monitoring sensor networks are platforms whose purpose is to monitor the evolution of certain phenomena in a specific area of interest or scenario, e.g., precision agriculture systems [18], air quality monitoring [19], or industrial processes [20, 21]. These networks are of particular interest because they monitor critical phenomena or assets and provide data for further applications or decision-making processes. Figure 1 depicts different scenarios where IoT monitoring networks are used, common use cases include smart metering, industrial applications, and environmental monitoring among others.
@@
Data-driven IoT applications aim to exploit the network measurements as well as their intrinsic structure to carry out several tasks. The quality of data from monitoring sensor networks is critical since the data can feed various applications or decision-making processes that require complete and accurate data. In the literature, data quality tasks have been developed to provide continuous, complete, and accurate data. Data quality in IoT platforms is defined by different dimensions such as data accuracy, sample completeness, data consistency or coherence, and data timeliness [22, 23, 24]. Some of the issues that may affect such data attributes are the presence of missing data, outliers, inconsistent samples, etc. Therefore, some of the tasks developed include the detection of outliers and anomalies [25], the use of predictive models to create virtual sensors [26], the imputation of missing data [27], and the detection of clusters and communities among others [16]. In the context of sensor data, models have also been developed to estimate the reliability of the measurements, an aspect related to possible measurement deviations and thus outlying or anomalous measurements [28]. Indeed, graph-based models provide alternatives to classic data-driven approaches and have been proven to be superior in terms of performance, possibility of distributed approaches, and interpretability. IoT data is characterized by the presence of spatiotemporal correlations given that monitoring platforms measure phenomena in an area of interest or assets in an industrial environment. In particular, graph-based models are known for their ability to model complex relationships between sensors and leverage spatiotemporal correlations [17]. Henceforth, a deep understanding of signal processing and ML over graph approaches may unveil new applications and advantages for IoT monitoring sensor networks.
@@
Recent surveys are based on the technical review of models based mainly on GSP and GNN, mentioning some applications in areas such as computer vision or natural language processing [14, 29, 30, 31, 32]. Others focus on the applications of GNNs and graph-based models in areas such as the IoT, showing graph-based applications in cases such as environmental monitoring or biological data [17, 33, 34, 35]. In this survey, we focus on the application of graph-based models to the recently booming area of data quality improvement for monitoring IoT platforms [36, 35]. In particular, we not only cover GNNs applications but also provide complete coverage of graph-based modeling of monitoring networks, covering GSP, ML over graphs, and GNNs. We also present recent advances and applications in data quality control and provide a taxonomy for the most novel graph-based applications and trends.
@@
The outline of this paper is as follows; Section 2 motivates the content of the paper. Section 3 introduces technical concepts regarding signal processing and ML over graphs. Section 4 reviews different graph applications in IoT monitoring sensor networks, and section 5 presents challenges, limitations, and emerging trends. Finally, section 6 concludes the manuscript. Figure 2 shows the outline of the survey.
@@ 2 Background and Motivation
Before the appearance of the GSP, graphs were already used in the field of graph-based semi-supervised learning to present the relationships of data residing in a manifold [29, 37]. In fact, label propagation emerged as a transductive method to estimate unlabeled data points from labeled ones, and inspired the GNN framework [38]. The advent of the GSP was a revolution since it allowed the translation of solid signal processing fundamentals to the realm of signals defined over graphs. Thus, this allowed the analysis of measurements residing on a graph via graph filters, signal shifts, signal convolutions, or the spectral graph Fourier transform among others. In turn, the graph signal convolution led to the translation of convolutional neural networks to the graph paradigm, making use of the graph signal convolution operator, e.g., the Chebnet GNN [39]. Besides all the guarantees in terms of efficiency and accuracy, the GSP-based techniques provide interpretability, which has received a lot of attention due to the growing use of black-box models, given the use of the graph shift operator, i.e., the matrix that describes the relationships between network nodes. Moreover, many of the techniques, e.g., graph filters, allow a distributed implementation, which is very interesting in case of resource limitations or in case of having nodes with computational resources, e.g., edge computing approach [40]. All this together makes the GSP provide appropriate tools to analyze the measurements produced by a network of sensors, being able to define these measurements as signals that reside in a graph.
@@
ML over graphs usually refers to ML techniques that have been adapted to take into account that a graph can represent the data, or analogously that the data resides in a manifold. In this sense, graph-based semi-supervised learning is aimed at the application of algorithms to partially-labeled data residing on manifolds represented by graphs [7]. This idea has also been used in different works where the optimization function of different ML techniques has been modified to include a regularization term that takes into account that the data points reside on a graph [41]. Among some examples, we can find predictive models, e.g., kernel ridge regression [42] or regularized support vector regression [43], which can be very useful in the case of sensor network measurements. Even though, there exists a large range of tasks and applications that can be approached using ML, e.g., the target of the ML task can be estimating features at graph nodes, at graph edges, or at the graph as a whole. Throughout the manuscript, we restrict to the tasks that are useful in the sensor network domain.
@@
Although GNNs belong to the field of ML over graphs, in this paper we want to dedicate their own space to them due to their great importance and widespread adoption nowadays. Some well-known architectures include the GraphSAGE [44] or the Chebnet [39]. Indeed, the graph convolution operator allowed the application of the convolution neural networks to the graph realm, resulting in graph convolutional neural networks such as the Chebnet, in which the convolution is realized via a graph filter. Given the widespread application of neural networks, tailored GNN architectures are being continuously developed to tackle tasks such as missing value imputation, anomaly detection, data compression, event detection, or signal reconstruction [14]. Another example of disruption in this field is the emergence of graph attention networks and graph transformers, being the analogs of attention mechanisms and transformers in conventional neural networks [45].
@@
Figure 3 depicts the scope of this review. From now on we focus on GSP, ML over graphs, and GNNs, placing special emphasis on novel applications for data quality in monitoring sensor networks. Data quality is a key challenge in monitoring sensor networks, with tasks ranging from anomaly detection to missing value imputation [36]. Although existing surveys cover various topics related to GSP and GNN for IoT (Table 1), the review of graph-based models for data quality in IoT has not yet been covered. In addition, the survey groups into useful applications (in the context of IoT data quality) different works that have developed methods for different applications in isolation. Hence, this survey provides a novel description of graph-based models for data quality tasks in IoT monitoring sensor networks. In summary, our contributions are:
@@
    describe the foundations of different graph-based models for IoT sensor networks (GSP, ML over graphs, and GNNs) that can be used in data quality enhancement applications,
@@
    review recent work and provide a taxonomy of applications that can help practitioners implement graph-based models in monitoring sensor networks to improve data quality and ensure data quality requirements,
@@
    highlight the challenges and limitations that practitioners may face, as well as identify emerging trends.
Table 1: A comparison between the scope of this survey and the scope of other relevant surveys.
@@ 3 Graph-Based Modeling for Sensor Networks
In this section, we elaborate on the fundamentals of the GSP, ML over graphs, and GNNs, highlighting current trends and novel techniques. We denote the measurement of a given sensor $i$ at a given time $t$ as $x_{i}(t){\in}\mathbb{R}$, $N$ denotes the number of sensors, and $\mathbf{x}_{i}{\in}\mathbb{R}^{T}$ represents a vector of $T$ measurements recorded by the $i$-th sensor. Matrices are denoted by bold uppercase letters, vectors by bold lowercase letters, and sets by calligraphic letters.
@@
Figure 4: Example of how a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{S}\}$ can be used to represent a sensor network and the sensors’ relationships. $e_{ij}$ represents an edge connecting nodes $i$ and $j$. The entry of the graph shift matrix $S_{ij}$ assigns a weight to the edge $e_{ij}$.
@@
Figure 5: Types of graphs: above, static graphics; below, dynamic graphics; $u$ denotes undirected, $d$ denotes directed, $m$ denotes multigraph, and $\mathcal{G}_{u}^{(t)}$ denotes a dynamic undirected graph at time $t$. $T{\in}\mathbb{N}$ represents a time offset.
In addition, there are two variants of graphs that can be applied in sensor networks: i) static graphs $\mathcal{G}$ and ii) dynamic or time-varying graphs $\mathcal{G}^{(t)}$ whose structure depends on time $t$ [55]. Static graphs can be used to represent sensor networks that are static or fixed so the relationships between the sensors are not expected to change. In contrast, dynamics graphs have a time-dependent structure, meaning that the relationships between the sensors as well as the set of sensors composing the graph can change over time. This approach is useful to represent mobile sensor networks or dynamic networks where sensors can join and leave the network and the relationship between them can change. We can further classify the graphs into three types; i) undirected, in which the relationships between nodes are reciprocal, i.e., $(i,j){\in}\mathcal{E}$ implies $(j,i){\in}\mathcal{E}$, or $S_{ij}{=}S_{ji}$, ii) directed, in which the relationships do not need to be reciprocal, and iii) multigraphs, in which more than one edge can exist for a given pair of nodes $(i,j)$. Figure 5 represents the different types of graphs. In the specific case of monitoring sensor networks, where a graph structure is used to represent the relationship between the sensors’ data, it makes sense to assume that the relationships between the sensor measurements are reciprocal, e.g., as in the case of correlation coefficients. This survey therefore focuses on the use of undirected static graphs, as monitoring platforms are typically fixed after deployment. More complex scenarios, such as mobile monitoring sensor networks, require more complex techniques and the use of dynamic graphs [^2] Refer to [56] for a systematic review of dynamic graphs..
@@
In addition, before moving on to explain the technical aspects of signal processing and ML over graphs, it is important to highlight the use of two types of predictive models in this graph-based approach; i) inductive models, and ii) transductive models. The objective of inductive learning is to infer a model from training instances, then this model can be applied to different testing instances. In contrast, in the case of transductive learning, training instances are used to obtain estimates for specific test instances. As we will see later, even though in the realm of sensor networks many models can be used in both ways, there are specific models of both types, such as a graph filter (inductive) or a graph kernel ridge regression (transductive). Figure 6 illustrates the idea behind these two learning paradigms.
@@
Figure 6: Inductive models are trained to perform inferences in other test sets while transductive models are trained to perform inferences for a specific test set.In black training data, and in blue testing data.